# Deciding Subspace reachability problems with application to Skolem’s Problem

### v1

@@ 2. Preliminaries
Let $\lambda_{1},...,\lambda_{m}$ label the eigenvalues of $A\in\mathbb{Q}^{d\times d}$, $m\leq d$. We always assume the roots are labelled so that $|\lambda_{1}|\geq|\lambda_{2}|\geq\cdots\geq|\lambda_{m}|$. In the case $|\lambda_{1}|=\cdots=|\lambda_{r}|>|\lambda_{r+1}|$ we say that $A$ has exactly $r$ roots of maximal modulus, or $r$ dominating roots.
In this paper we work in ambient Euclidean space $\mathbb{R}^{d}$, with orbits $\{A^{n}x\}_{n=0}^{\infty}$ contained in $\mathbb{Q}^{d}$. All standard algebraic operations, such as sums, products, root finding of polynomials, and computing Jordan canonical forms of rational matrices [cai1994computing] are well-known to be computable, and hence we do not discuss the details of such procedures (see Section 2.1 for details on the effective manipulation of algebraic numbers).
@@
For any matrix $A\in\mathbb{R}^{d\times d}$, there exists a basis in which the matrix of $A$ is a quasi-diagonal matrix $J=\text{Diag}(J_{1},...,J_{N})$ where each block $J_{i}$ is of form
@@
where the $\lambda_{t}$, $t=1,...,u$ are the real eigenvalues of $A$, and the $\lambda_{l},\bar{\lambda}_{l}=\alpha_{l}\pm i\beta_{l}$, $l=1,...,s$, $\beta>0$ are the complex eigenvalues of $A$. The sizes of the blocks are determined by the elementary divisors of $A$.
As such, we have $A=QJQ^{-1}$, and if $A$ is an algebraic matrix, then $J$ and $Q$ are also algebraic matrices and their entries can be computed from the entries of $A$. The usefulness of the real Jordan canonical form follows from the fact that in the case the characteristic polynomial of $A$ has complex roots, the Jordan blocks of $J$ continue to have real entries. For further details on the real Jordan canonical form including proof of Lemma 1 and effective methods for computing $J$, consult [shilov2012linear]. For the remainder of this paper, we suppose $J$ is the real Jordan canonical form of $A$. In addition, suppose Jordan blocks $J_{i}$, $i=1,...,N$ are of dimension $D_{i}$, and associated with eigenvalues $\lambda_{i}$. Of course, there may be many Jordan blocks associated with a single eigenvalue $\lambda_{i}$, so $N\geq m$, but we often abuse notation and say $J_{i}$ associates to $\lambda_{i}$ to speak of the collection of Jordan blocks associated with $\lambda_{i}$ — in the event more precise language is needed we use it.
@@
Letting $Q^{-1}x=x^{\prime}$ and $Q^{-1}U=U^{\prime}$, it follows that $A^{k}x\in U$ if and only if $Jx^{\prime}\in U^{\prime}$. ∎
@@ 2.1. Algebraic numbers and computing eigenvalues, eigenvectors
where $d=deg(p)$ and $H=H(p)$. As a consequence, when $r$ is less than the root separation bound, we have equality checking between algebraic numbers. Given distinct $\alpha,\beta\in\mathbb{A}$ these are roots of $p_{\alpha}p_{\beta}$, which is of degree at most $deg(\alpha)+deg(\beta)$, and of height at most $H(\alpha)H(\beta)$. Then one may compute $\alpha+\beta$, $\alpha\beta$, $1/\alpha$, $\overline{\alpha}$, $|\alpha|$, decide whether $\alpha>\beta$ for distinct algebraic $\alpha$ and $\beta$, and more.
@@ 2.2. Linear recurrence sequences
where $a_{0},...,a_{d-1}\in\mathbb{Q}$, with $a_{0}\neq 0$ and $u_{j}\neq 0$ for at least one $j$ in the range $10\leq j\leq m-$ 1. We say such an LRS has order $d$. We call the $a_{0},...,a_{d-1}$ the coefficients of the sequence $\{u_{n}\}$ and the initial terms of $\{u_{n}\}$ are $u_{0},...,u_{d-1}$. The characteristic polynomial of the sequence $\{u_{n}\}$ is
@@
with the $\lambda_{1},...,\lambda_{k}$ called the characteristic roots of the sequence $\{u_{n}\}$.
@@
to be the companion matrix of the LRS, and take the vector $x\in\mathbb{Q}^{d}$ as $(u_{d-1},...,u_{0})^{T}$ of initial terms of the LRS. Then iteration of $A$ over $x$ acts as a “shift” on the entries of $x$, shifting in the next term of the LRS and dropping the oldest term. In addition, the characteristic polynomial of the LRS is the characteristic polynomial of $A$, and the characteristic roots of the LRS are the eigenvalues of $A$.
@@ 2.3. Angles between flats
Let $U,W\subset\mathbb{R}^{d}$ be $k$ and $l$ dimensional subspaces, respectively, presented by orthogonal rational basis $\{u_{1},...,u_{k}\}$ and $\{w_{1},...,w_{l}\}$. Note any linearly independent set of rational vectors can be effectively transformed into an orthogonal set of rational vectors via the Gram-Schmidt Process (before normalization). If instead it is desired that the basis be orthonormal, then the entries of the vectors will be algebraic after the renormalization step.
In this paper we are primarily concerned with determining the minimal angle $20\leq\theta_{1}\leq\pi/$ 2 between two subspaces, defined as the least angle between any pair of non-zero vectors from $U$ and $W$. This is the first principle angle between subspaces, whose variational characterization is
(1) $\theta=\min\left\{\arccos\left(\frac{|\langle u,w\rangle|}{||u||~{}||w||}\right):u\in U,w\in W\right\},$
[…]
@@
$\sigma_{1}=\max_{x\in\mathbb{C}^{m},y\in\mathbb{C}^{n}}\frac{|x^{*}Ay|}{||x||\ ||y||},$
where $\sigma_{1}$ is the first singular value of $A$. Dropping the $\arccos$ term in Equation 1, we obtain a new problem, namely maximizing the quantity
(2) $\sigma=\max\left\{\frac{|\langle u,w\rangle|}{||u||\ ||w||}:u\in U,w\in W\right\}.$
@@
Combining the above observations, take matrices $X\in\mathbb{R}^{n\times k}$ and $Y\in\mathbb{R}^{n\times l}$ to have columns consisting of orthonormal bases $\{u_{1},...,u_{k}\}$ and $\{w_{1},...,w_{l}\}$ of $U$ and $W$, respectively. The optimization problem in Equation 2 can then be written as
$\sigma=\max_{x\in\mathbb{R}^{k},y\in\mathbb{R}^{l}}\frac{x^{\top}X^{\top}Yy}{||x||\ ||y||},$
the solution to which is the largest singular value of $X^{\top}Y$, by the variational characterization of the singular value decomposition. For more details and proof of the fact that cosines of principle angles come from the singular value decomposition see [knyazev2002principal], or refer to Chapter 5 of [meyer2023matrix] for more details on the variational characterization of singular value decomposition, and additional effective methods for computing the minimal angle between subspaces.
@@
(3) $\Gamma(U,W)=\max_{x\in\mathbb{R}^{k},y\in\mathbb{R}^{l}}\frac{x^{\top}X^{\top}Yy}{||x||\ ||y||}=\sigma,$
to be the function computing the minimal angle between subspaces $U$ and $W$ as above. Whether $\Gamma(U,W)$ returns $\sigma$ or $\sigma^{2}$ does not matter for the purposes of this paper — intuitively $\Gamma(U,W)$ computes the minimal angle between the subspaces ($\arccos\Gamma$ is the actual minimal cosine angle), and everything remains computable: the columns of $U$ and $W$ will always be composed of rational (or possibly algebraic) numbers, and $\Gamma$ is effectively computable with such input.
2.4. Angle evolution, and the Orbit Problem in real projective space
Real projective space $\mathbb{R}\mathbb{P}^{d-1}$ is defined to be the quotient of $\mathbb{R}^{d}\setminus\{0\}$ by the equivalence relation $x\sim\alpha x$ where $\alpha\neq 0$ is real and $x\in\mathbb{R}^{d}$. Clearly, it suffices to consider only vectors $x$ with Euclidean norm $||x||=1$. Hence, geometrically, projective space is the space of lines through the origin, or alternatively the space obtained by identifying opposite points on the unit sphere $S^{d-1}$. The projective linear group $PGL(\mathbb{R}^{d})=GL(\mathbb{R}^{d})/\alpha\cdot\text{Id}$, $\alpha\neq 0$, is the induced action of the general linear group on projective space. The induced action of non-invertible square matrices over elements of projective space is similarly defined.
@@ 2.4. Angle evolution, and the Orbit Problem in real projective space
$d(\mathbb{P}x,\mathbb{P}y)=\min\left(\left|\left|\frac{x}{||x||}-\frac{y}{||y||}\right|\right|,\left|\left|\frac{x}{||x||}-\frac{-y}{||y||}\right|\right|\right).$
In the case that $W$ is a subspace of $\mathbb{R}^{d}$, define
$d(x,W\cap S^{d-1})=\inf_{y\in W\cap S^{d-1}}||x-y||=\min_{y\in W}d(\mathbb{P}x,\mathbb{P}y)=d(\mathbb{P}x,\mathbb{P}W).$
@@
$\Gamma(A^{n}x,U)=\frac{|A^{n}x\cdot y|}{||A^{n}x||\ ||y||},\ n\in\mathbb{N}.$
@@
$\frac{|A^{n}x\cdot y|}{||A^{n}x||\ ||y||}\rightarrow 1\text{ as }n\rightarrow\infty,$
@@
$d(\mathbb{P}A^{n}x,\mathbb{P}y)=\min\left(\left|\left|\frac{A^{n}x}{||A^{n}x||}-\frac{y}{||y||}\right|\right|,\left|\left|\frac{A^{n}x}{||A^{n}x||}-\frac{-y}{||y||}\right|\right|\right)\rightarrow 0\text{ as }n\rightarrow\infty.$
@@
$\Gamma(U,W)=\sigma=\min\left\{\left(\frac{|\langle u,w\rangle|}{||u||~{}||w||}\right):u\in U,w\in W\right\}.$
@@
$\Gamma(A^{n}x,W)=\sigma=\min\left\{\left(\frac{|\langle A^{n}x,w\rangle|}{||A^{n}x||~{}||w||}\right):w\in W\right\}.$
@@
$d(\mathbb{P}A^{n}x,\mathbb{P}W)=\inf_{y\in W\cap S^{d-1}}||A^{n}x-y||=\min_{y\in W}d(\mathbb{P}A^{n}x,\mathbb{P}y),$
@@
Given $x\in\mathbb{R}^{d}$, let $X$ denote the one-dimensional subspace (line) spanned by $x$. Then, for a matrix $A\in\mathbb{R}^{d\times d}$ and subspace $U$, we see that there exists $n\in\mathbb{N}$ such that $A^{n}x\in U$ if and only if $A^{n}X\subseteq U$, where $A^{n}X$ is the subspace spanned by the vector $A^{n}x$. Hence, given an instance $(A,x,U)$ of the Orbit Problem, it is immediate that there is an $n\in\mathbb{N}$ such that $A^{n}x\in U$ if and only if $\mathbb{P}A^{n}x\in\mathbb{P}U$.
@@
We begin by stating the following well-known result about Jordan blocks, easily proved by induction.

### v2

@@ 2. Preliminary results
Let $\lambda_{1},\dots,\lambda_{m}$ label the eigenvalues of $A\in\mathbb{Q}^{d\times d}$, $m\leq d$. We always assume the roots are labelled so that $|\lambda_{1}|\geq|\lambda_{2}|\geq\cdots\geq|\lambda_{m}|$. In the case $|\lambda_{1}|=\cdots=|\lambda_{r}|>|\lambda_{r+1}|$ we say that $A$ has $r$ roots of maximal modulus, or $r$ dominating roots.
In this paper we work in ambient Euclidean space $\mathbb{R}^{d}$ (before compactifying), with orbits $\{A^{n}x\}_{n=0}^{\infty}$ contained in $\mathbb{Q}^{d}$. All standard algebraic operations, such as sums, products, root finding of polynomials, and computing Jordan canonical forms of rational matrices [cai1994computing] are well-known to be computable, and hence we do not discuss the details of such procedures (see Section 2.1 for details on the effective manipulation of algebraic numbers).
@@
For any matrix $A\in\mathbb{R}^{d\times d}$, there exists a basis in which the matrix of $A$ is a quasi-diagonal matrix $J=\text{Diag}(J_{1},\dots,J_{N})$ where each block $J_{i}$ is of form
@@
where the $\lambda_{t}$, $t=1,\dots,u$ are the real eigenvalues of $A$, and the $\lambda_{l},\bar{\lambda}_{l}=\alpha_{l}\pm i\beta_{l}$, $l=1,\dots,s$, $\beta>0$ are the complex eigenvalues of $A$. The sizes of the blocks are determined by the elementary divisors of $A$.
As such, we have $A=QJQ^{-1}$, and if $A$ is algebraic (has algebraic entries), then $J$ and $Q$ are also algebraic matrices and their entries can be computed from the entries of $A$. The usefulness of the real Jordan canonical form follows from the fact that in the case the characteristic polynomial of $A$ has complex roots, the Jordan blocks of $J$ continue to have real entries. For further details on the real Jordan canonical form including proof of Lemma 1 and effective methods for computing $J$, consult [shilov2012linear].
For the remainder of this paper, we suppose $J$ is the real Jordan canonical form of $A$. In addition, suppose Jordan blocks $J_{i}$, $i=1,\dots,N$ are of dimension $D_{i}$, and associated with eigenvalues $\lambda_{i}$. Of course, there may be many Jordan blocks associated with a single eigenvalue $\lambda_{i}$, so $N\geq m$, but we often abuse notation and say $J_{i}$ associates to $\lambda_{i}$ to speak of the collection of Jordan blocks associated with $\lambda_{i}$ — in the event more precise language is needed we use it. Moreover, we remark that the largest Jordan blocks associated with an eigenvalue are determined by the multiplicity of the eigenvalue in the minimal polynomial of $J$ — this is a useful fact used throughout our paper.
@@
Letting $Q^{-1}x=x^{\prime}$ and $Q^{-1}U=U^{\prime}$, it follows that $A^{k}x\in U$ if and only if $J^{k}x^{\prime}\in U^{\prime}$. ∎
@@ 2.1. Algebraic numbers and computing eigenvalues, eigenvectors
where $d=deg(p)$ and $H=H(p)$. As a consequence, when $r$ is less than the root separation bound, we have equality checking between algebraic numbers. Given distinct $\alpha,\beta\in\mathbb{A}$ these are roots of $p_{\alpha}p_{\beta}$, which are of degree at most $deg(\alpha)+deg(\beta)$, and of height at most $H(\alpha)H(\beta)$. Then one may compute $\alpha+\beta$, $\alpha\beta$, $1/\alpha$, $\overline{\alpha}$, $|\alpha|$, decide whether $\alpha>\beta$ for distinct algebraic $\alpha$ and $\beta$, and more.
@@ 2.2. Linear recurrence sequences
where $a_{0},\dots,a_{d-1}\in\mathbb{Q}$, with $a_{0}\neq 0$ and $u_{j}\neq 0$ for at least one $j$ in the range $0\leq j\leq d-1$. We say such an LRS has order $d$. We call the $a_{0},\dots,a_{d-1}$ the coefficients of the sequence $\{u_{n}\}$ and the initial terms of $\{u_{n}\}$ are $u_{0},\dots,u_{d-1}$. The characteristic polynomial of the sequence $\{u_{n}\}$ is
@@
with the $\lambda_{1},\dots,\lambda_{k}$ called the characteristic roots of the sequence $\{u_{n}\}$.
@@
to be the companion matrix of the LRS, and take the vector $x\in\mathbb{Q}^{d}$ as $(u_{d-1},\dots,u_{0})^{T}$ of initial terms of the LRS. Then iteration of $A$ over $x$ acts as a “shift” on the entries of $x$, shifting in the next term of the LRS and dropping the oldest term. In addition, the characteristic polynomial of the LRS is the characteristic polynomial of $A$, and the characteristic roots of the LRS are the eigenvalues of $A$.
@@ 2.3. Angles between flats
Let $U,W\subset\mathbb{R}^{d}$ be $k$ and $l$ dimensional subspaces, respectively, presented by orthogonal rational basis $\{u_{1},\dots,u_{k}\}$ and $\{w_{1},\dots,w_{l}\}$. Note any linearly independent set of rational vectors can be effectively transformed into an orthogonal set of rational vectors via the Gram-Schmidt Process (before normalization). If instead it is desired that the basis be orthonormal, then the entries of the vectors will be algebraic after the renormalization step.
In this paper we are primarily concerned with determining the minimal angle $0\leq\theta_{1}\leq\pi/2$ between two subspaces, defined as the least angle between any pair of non-zero vectors from $U$ and $W$. This is the first principal angle between subspaces, whose variational characterization is
(1) $\theta=\min\left\{\arccos\left(\frac{|\langle u,w\rangle|}{\|u\|~\|w\|}\right):u\in U,w\in W,\text{ and }u,w\neq 0\right\},$
where $\langle\cdot,\cdot\rangle$ is the standard inner (dot) product. We refer to the quantity expressed above in Equation 1 as the minimal angle between subspaces $U$ and $W$. We always assume $U$ and $W$ have trivial intersection $U\cap W=\{0\}$, so that the minimal angle is well-defined and non-zero.
@@
$\sigma_{1}=\max_{x\in\mathbb{C}^{m},y\in\mathbb{C}^{n}}\frac{|x^{*}Ay|}{\|x\|\ \|y\|},$
where $x^{*}$ is the conjugate transpose and $\sigma_{1}$ is the first singular value of $A$. Dropping the $\arccos$ term in Equation 1, we obtain a new problem, namely maximizing the quantity
(2) $\sigma=\max\left\{\frac{|\langle u,w\rangle|}{\|u\|\ \|w\|}:u\in U,w\in W\right\}.$
@@
Combining the above observations and restricting to the reals, take matrices $X\in\mathbb{R}^{n\times k}$ and $Y\in\mathbb{R}^{n\times l}$ to have columns consisting of orthonormal bases $\{u_{1},\dots,u_{k}\}$ and $\{w_{1},\dots,w_{l}\}$ of $U$ and $W$, respectively. The optimization problem in Equation 2 can then be written as
$\sigma=\max_{x\in\mathbb{R}^{k},y\in\mathbb{R}^{l}}\frac{x^{\top}X^{\top}Yy}{\|x\|\ \|y\|},$
the solution to which is the largest singular value of $X^{\top}Y$, by the variational characterization of the singular value decomposition. For more details and proof of the fact that cosines of principal angles come from the singular value decomposition see [knyazev2002principal], or refer to Chapter 5 of [meyer2023matrix] for more details on the variational characterization of singular value decomposition, and additional effective methods for computing the minimal angle between subspaces.
@@
(3) $\Gamma(U,W)=\max_{x\in\mathbb{R}^{k},y\in\mathbb{R}^{l}}\frac{x^{\top}X^{\top}Yy}{\|x\|\ \|y\|}=\sigma,$
to be the function computing the minimal angle between subspaces $U$ and $W$ as above. If one or both of $U$ and $W$ are one-dimensional subspaces, i.e. $U=\text{span}(x)$ for $x\in\mathbb{Q}^{d}$, then $\Gamma(x,W)$ is understood to mean $\Gamma(U,W)$ to allow for more natural notation.
Whether $\Gamma(U,W)$ returns $\sigma$ or $\sigma^{2}$ does not matter for the purposes of this paper — intuitively $\Gamma(U,W)$ computes the minimal angle between the subspaces ($\arccos\Gamma$ is the actual minimal cosine angle), and everything remains computable: the columns of $U$ and $W$ will always be composed of rational (or possibly algebraic) numbers, and $\Gamma$ is effectively computable with such input.
2.4. The Orbit Problem in real projective space
Real projective space $\mathbb{R}\mathbb{P}^{d-1}$ is defined to be the quotient of $\mathbb{R}^{d}\setminus\{0\}$ by the equivalence relation $x\sim\alpha x$ where $\alpha\neq 0$ is real and $x\in\mathbb{R}^{d}$. Clearly, it suffices to consider only vectors $x$ with Euclidean norm $\|x\|=1$. Hence, geometrically, projective space is the space of lines through the origin, or alternatively the space obtained by identifying opposite points on the unit sphere $S^{d-1}$. The projective linear group $PGL(\mathbb{R}^{d})=GL(\mathbb{R}^{d})/\alpha\cdot\text{Id}$, $\alpha\neq 0$, is the induced action of the general linear group on projective space. The induced action of non-invertible square matrices over elements of projective space is similarly defined.
@@ 2.4. The Orbit Problem in real projective space
[…]
In the case that $W$ is a subspace of $\mathbb{R}^{d}$, define the distance between a point $x$ and the subspace to be
$d(x,W\cap S^{d-1})=\inf_{y\in W\cap S^{d-1}}\|x-y\|=\min_{y\in W}d(\mathbb{P}x,\mathbb{P}y)=d(\mathbb{P}x,\mathbb{P}W).$
@@
$\Gamma(A^{n}x,U)=\frac{|A^{n}x\cdot y|}{\|A^{n}x\|\ \|y\|},\ n\in\mathbb{N}.$
@@
$\frac{|A^{n}x\cdot y|}{\|A^{n}x\|\ \|y\|}\rightarrow 1\text{ as }n\rightarrow\infty,$
@@
[…]
@@
$\Gamma(U,W)=\sigma=\max\left\{\left(\frac{|\langle u,w\rangle|}{\|u\|~\|w\|}\right):u\in U,w\in W\right\}.$
@@
$\Gamma(A^{n}x,W)=\sigma=\max\left\{\left(\frac{|\langle A^{n}x,w\rangle|}{\|A^{n}x\|~\|w\|}\right):w\in W\right\}.$
@@
$d(\mathbb{P}A^{n}x,\mathbb{P}W)=\inf_{y\in W\cap S^{d-1}}\|A^{n}x-y\|=\min_{y\in W}d(\mathbb{P}A^{n}x,\mathbb{P}y),$
@@
Given $x\in\mathbb{R}^{d}$, let $X$ denote the one-dimensional subspace (line) spanned by $x$. Then, for a matrix $A\in\mathbb{R}^{d\times d}$ and subspace $U$, we see that for every $n\in\mathbb{N}$, $A^{n}x\in U$ if and only if $A^{n}X\subseteq U$, where $A^{n}X$ is the subspace spanned by the vector $A^{n}x$. Hence, given an instance $(A,x,U)$ of the Orbit Problem, it is immediate that there is an $n\in\mathbb{N}$ such that $A^{n}x\in U$ if and only if $\mathbb{P}A^{n}x\in\mathbb{P}U$.
@@
2.5. Convergence of orbits toward attracting subspaces in $\mathbb{RP}^{d-1}$
In this section we establish the lemmas underlying Theorem 1, and hence all the other results in this paper. The following lemmas and the arguments used are technical, which perhaps obfuscates the underlying intuition and central concepts which are in-fact quite simple. To this end, we take a moment to present the intuition behind the following results, which should aid in clarifying one’s understanding.
The central concept is this. Let
$\Lambda=\text{Diag}(x_{1},x_{2},\dots,x_{d})$
be a diagonal matrix with $x_{1}>x_{2}>\cdots>x_{d}\geq 1$. If $y\in\mathbb{Q}^{d}$, then iteration of $\Lambda$ over $y$ sends the first entry of $y$ to $\infty$ faster than the others since $x_{1}>x_{i}$, $i\neq 1$, i.e.
$\lim_{n\rightarrow\infty}\frac{x_{i}^{n}}{x_{1}^{n}}=0\text{ for }i\neq 1.$
Hence, $\Gamma(\Lambda^{n}y,e_{1})\rightarrow 1$ as $n\rightarrow\infty$ where $e_{1}$ is the first standard basis vector. In particular, following Lemma 3,
$\lim_{n\rightarrow\infty}d(\mathbb{P}\Lambda^{n}x,\mathbb{P}e_{1})=0.$
This observation underlies the idea behind Theorem 1: if the orbit of a point in projective space approaches an attracting subspace that does not intersect with the target subspace, then eventually the orbit must be bound away form the target subspace for all time, so we have only a finite number of iterations for which intersection must be checked.
The following lemmas deal with a generalization of the above example, where $\Lambda$ is not a real diagonal matrix but rather a matrix in Jordan normal form. However, one may just as easily abstractly treat the $x_{i}$ composing $\Lambda$ not as real numbers, but rather as Jordan blocks. Then the same principle holds: a certain Jordan block or collection of identical Jordan blocks will “dominate” the asymptotics, and hence $\Gamma(\Lambda^{n}y,E)\rightarrow 1$ as $n\rightarrow\infty$. The critical difference being that $E$ may not be one-dimensional, but is a direct sum of the one or two-dimensional subspaces invariant under each real or complex Jordan block dominating the asymptotics, respectively.
Equipped with this intuition, we begin by stating the following well-known result about Jordan blocks, easily proved by induction.